OpenStudy (anonymous):

- URGENT- Today I spent half the money I had. The amount remaining in 'pounds and pence' is the same as the original amount in 'pence and pounds'. How much did I have to start with? SHOW WORKING OUT/LOGIC PLEASE 1 pound = 100 pence

OpenStudy (anonymous):

pounds

jhonyy9 (jhonyy9):

1 pound = ? pences

jhonyy9 (jhonyy9):

the amount remaining in POUNDS and PENCE is the same as the original amount in PENCE and POUNDS so ATTENTION the first amount of POUNDS is the same with remained amount of PENCE and viceversa the remained amount of POUNDS is the same with first amount of PENCE - i think this will be the key to solve right this exercise

jhonyy9 (jhonyy9):

smith what is your opinion about please ?

OpenStudy (agent0smith):

I've tried solving it, but it seems pretty tricky. If you started with x pounds and y pence, then you must have ended with y pounds and x pence. And you spent half of the initial amount. But I haven't really got anywhere with that.

jhonyy9 (jhonyy9):

ok. but how you think it now this make sens how i have thought it so in this way ? thank you

OpenStudy (agent0smith):

Yeah I agree with you.

jhonyy9 (jhonyy9):

yes but you know again that 1 pound equal 100 pence suppose start with 80 pound 10 pence at the end remained 80 pence 10 pounds the difference 80 - 10 =70 pounds and 10-80=-70 so will be spent 69 pounds 30 pences - hope i have calculed it right now with this convert

jhonyy9 (jhonyy9):

smith - thank you for you agree this way and hope so much that now this is understandably sure right easy - yes ?

jhonyy9 (jhonyy9):

ohhhh ok yes you are right

OpenStudy (agent0smith):

It doesn't appear to be easy @jhonyy9

jhonyy9 (jhonyy9):

yes now i see it - but the way need being this i have wrote above - yes ?

jhonyy9 (jhonyy9):

so 1. start with x pounds and y pences and will remained x/2 pounds and y/2 pences x=y/2 =>2x=y halfed (2x+y)/2 = y=x/2 => 2y=x smith opinion about ?

OpenStudy (agent0smith):

I think I got it. I'm not going to show all the working for solving the equation, since I'm gonna leave that part up to you. I'm not here to do boring algebra. We start with x pounds and y pence, so our starting money value is: x + 0.01y We finish with y pounds and x pence, so our final money value is y + 0.01x. You spent half the starting amount, so the final amount is equal to half the starting amount: \[\large \frac{ 1 }{ 2 }(x+0.01y) = y +0.01x\]

OpenStudy (agent0smith):

http://www.wolframalpha.com/input/?i=(x%2B0.01y)%2F2+%3D+y%2B0.01x+integer+solutions Simplest solution is starting with x = 199 pounds and y = 98 pence, total value is 199.98 After spending half that amount, you'd have 98 pounds and 199 pence... total value of 99.99.

jhonyy9 (jhonyy9):

yes i think make sens right but than i see it right at the start the pence part is or wan being x/100 - yes ?

jhonyy9 (jhonyy9):

@agent0smith very very nice job - work - Congratulation @agent0smith !!!

OpenStudy (agent0smith):

lol I don't care about medals. I'm not here to do every single step of working. You should be doing (or at least attempting to do) some of the work. I just took it as a challenge to solve this.

jhonyy9 (jhonyy9):

smith - can you please make the Asker understanding the right way how he or she can solving this exercise - please because i think you can make it more easy than i - thank you in advance

jhonyy9 (jhonyy9):

@agent0smith - please

OpenStudy (anonymous):

1 pound

OpenStudy (anonymous):

because the amount you got was some pound

OpenStudy (agent0smith):

Sorry @jhonyy9 I had to leave earlier.